OpenStudy (anonymous):

This is a Quadratic Equation, solve it by factoring please. -x^2+6x-5=0

OpenStudy (anonymous):

and please explain

OpenStudy (anonymous):

-(x-1)(x-5)

OpenStudy (anonymous):

First get it rid of the - in front of the x^2 by multiplying by -1

OpenStudy (anonymous):

You get \[ x^2-(1+5)x+1*5 = (x-1)(x-5) = 0 \] So the solutions are 1 and 5.

OpenStudy (anonymous):

thank youu